if Don is 21 years older than betty, and in 6 years he will twice as old as betty how old is each person
Answer by scott8148(6628)

d = b + 21

d + 6 = 2(b + 6)

substituting ___ (b + 21) + 6 = 2(b + 6) ___ b + 27 = 2b + 12

solve for b, then substitute back to find d